OpenStudy (anonymous):

Write the polar form of the complex number 3-2i, using θ such that 0 ≤ θ ≤ 360º. Round numerical values to the nearest hundredth.

OpenStudy (helder_edwin):

if u have \[ z=a+ib \] then its polar form is \[\large z=\rho(\cos\theta+i\sin\theta) \]